OpenStudy (anonymous):

What is the simplified form of 4x^2-25/2x-5?

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

hint: we can factorize the numerator as follows: \[\Large \frac{{4{x^2} - 25}}{{2x - 5}} = \frac{{\left( {2x - 5} \right)\left( {2x + 5} \right)}}{{2x - 5}} = ...?\]

OpenStudy (anonymous):

2x+5

OpenStudy (michele_laino):

since in general we have: \[\Large {a^2}{x^2} - {b^2} = \left( {ax - b} \right)\left( {ax + b} \right)\]

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

What would the restriction be?

OpenStudy (michele_laino):

since we can not divide by zero, then we have to exclude those values of x such that: \[\Large 2x - 5 = 0\]

OpenStudy (anonymous):

So -5/2

OpenStudy (michele_laino):

we have to exclòude x=5/2

OpenStudy (michele_laino):

exclude*

OpenStudy (anonymous):

So I was wrong when I said -5/2?

OpenStudy (michele_laino):

yes! since \[2x - 5 = 0\] when: x=5/2

OpenStudy (michele_laino):

please try to solve that equation: 2x-5=0

OpenStudy (anonymous):

Ok! I see! So the answer would be 2x+5, with a restriction of 5/2

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

:)!!! May I ask another?

OpenStudy (michele_laino):

ok!

OpenStudy (anonymous):

What is the simplifed form of